### v2

# Semantic Density: Uncertainty Quantification in Semantic Space for Large Language Models

Abstract
With the widespread application of Large Language Models (LLMs) to various domains, concerns regarding the trustworthiness of LLMs in safety-critical scenarios have been raised, due to their unpredictable tendency to hallucinate and generate misinformation. Existing LLMs do not have an inherent functionality to provide the users with an uncertainty metric for each response it generates, making it difficult to evaluate trustworthiness. Although a number of works aim to develop uncertainty quantification methods for LLMs, they have fundamental limitations, such as being restricted to classification tasks, requiring additional training and data, considering only lexical instead of semantic information, and being prompt-wise but not response-wise. A new framework is proposed in this paper to address these issues. Semantic density extracts uncertainty information for each response from a probability distribution perspective in semantic space. It has no restriction on task types and is "off-the-shelf" for new models and tasks. Experiments on seven state-of-the-art LLMs, including the latest Llama 3 and Mixtral-8x22B models, on four free-form question-answering benchmarks demonstrate the superior performance and robustness of semantic density compared to prior approaches.
1 Introduction
Large language models (LLMs) have revolutionalized many domains, such as conversational agents [27], code generation [33], and mathematical discovery [31]. Given their ability for general reasoning and adaptability to new tasks, LLMs are utilized increasingly in safety-critical applications, including healthcare [34] and finance [42]. However, existing LLMs have an unpredictable tendency to hallucinate [24], leading to misleading information and risky behaviors. Responses are generated without quantitative indicators for their uncertainty, making it difficult to evaluate how trustworthy they are. As a result, concerns have been raised about their safety [46], hindering a deeper utilization of LLMs in risk-sensitive domains [4].
Although significant resources have been invested in LLM development, leading to a rapid pace in new model releases, only little progress has been made in building an uncertainty quantification framework for LLMs. An ideal outcome of such a system would be a quantitative metric associated with each response that can be used as an uncertainty/confidence indicator. Users can then build on this metric to evaluate the trustworthiness of LLM responses, e.g., establish an automatic system that triggers a warning if the response uncertainty reaches a pre-defined threshold.
Following this line of thought, several techniques have been proposed in the literature to extract the uncertainty/confidence score from LLMs. In addition to the baselines that directly ask the LLM itself to evaluate its own answers [36, 15], one further step was to integrate traditional uncertainty estimatition/calibration methods into LLMs [5, 43, 45]. However, due to the nature of these traditional methods, they only work on classification problems, not free-form natural language generation (NLG) tasks, which are more general and challenging. Another direction was to fine-tune the original model [19] or train an additional layer or classifier [3, 22] to output uncertainty indicators for the responses. The main drawback is that these approaches are not "off-the-shelf" for new tasks and models: additional task-specific training labels of the ground-truth confidence are needed, and the training needs to be done in a model-specific manner, limiting their applicability.
Most importantly, prior work still treats LLM outputs as traditional auto-regressive predictions [23], i.e., the generated responses are simply handled as sequences of tokens/words, considering only their lexical uncertainty. However, due to the unique nature of free-form NLG, tokens that are lexically different may be semantically similar. In most LLM applications, decisions depend on the semantics of the generated responses, and the same semantics can be stated using different words or sentence structures, leading to different lexical tokens. Therefore, uncertainty in semantic space is a more essential indicator for trustworthiness of LLM responses than lexical uncertainty.
Semantic entropy [16] is the state-of-the-art (SOTA) technique in semantic uncertainty [21]. However, its current design has two intrinsic limitations. First, the returned uncertainty score is prompt-wise, i.e., the semantic entropy is calculated for each prompt, instead of each response. Considering that LLMs can generate diverse responses for the same prompt, using the same uncertainty score for different responses is problematic [20]. Second, semantic entropy only evaluates semantic equivalence between the generated responses, which is a binary one-cut measurement that returns whether two responses are semantically equivalent or not. It doesn’t take into account the more fine-grained semantic differences between the responses, which encode information that can make uncertainty quantification more precise.
To fill these gaps, a framework is developed in this paper for a new uncertainty metric, semantic density (SD), that can quantify the uncertainty of LLM responses in semantic space. Semantic density rebuilds the output probability distribution from a semantic perspective, and extracts an uncertainty indicator analogous to probability density. The proposed semantic density metric has the following advantages: (1) It does not need any further training or fine-tuning of the original LLM; it is an "off-the-shelf" tool that can be directly applied to any pre-trained LLMs without modifying them; (2) it does not pose any restrictions on the problem type; in particular, it works for general free-form generation tasks (3) the returned metric is response-wise, making it possible to evaluate trustworthiness of each response; and (4) it takes the fine-grained semantic differences between resposes into account, which makes uncertainty quantification more precise.
The performance of the semantic density metric was compared with six existing uncertainty quantification methods designed for LLMs across four question-answering benchmark datasets. All the approaches were tested on seven SOTA LLMs, including the latest Llama 3 and Mixtral-8x22B models. Semantic density performed significantly better than the alternatives across the board, suggesting that it forms a promising foundation for evaluating trustworthiness of LLM responses.
2 Related Work
Although the main focus of LLM community is still on developing new models with better performance, a number of studies aim at measuring uncertainty in LLMs. This section summarizes their basic ideas and potential limitations, which are then targeted in the development of semantic density.
The first direction is to ask the LLM to evaluate the uncertainty of its own responses. Tian et al. [36] performed an empirical study showing that the inherent conditional probabilities are poorly calibrated in existing LLMs with RLHF (reinforcement learning from human feedback), and that verbal confidence estimates provided by the LLMs are better calibrated. Kadavath et al. [15] developed an approach where the LLM was asked to evaluate the correctness of its own answer, in the form of the probability "P(True)" that its own answer is correct. An additional “value” head was also trained to predict "P(True)", but it turned out not to generalize well to out-of-distribution datasets. In general, the performance of model self-evaluation is not as good as other more advanced uncertainty quantification methods [16].
A second direction is to integrate traditional uncertainty quantification methods into LLMs. The effectiveness of temperature scaling [8] for calibrating the output token probabilities of LLMs was verified in Desai and Durrett [5] and Xiao et al. [43]. Ye et al. [45] utilize conformal prediction to quantify the uncertainty of LLMs. However, these approaches are limited to NLP classification tasks; in contrast, the proposed semantic density can be applied to the more general and challenging free-form generation tasks.
A third direction is to perform supervised learning, either by fine-tuning the original LLM or adding additional layer/classifier, to create uncertainty indicators. For example Lin et al. [19] fine-tuned the GPT-3 model to verbally express its own confidence level, and Azaria and Mitchell [3] trained an additional classifier to return the truthful probability of each generated response, based on the hidden layer activations of the LLM. Liu et al. [22] proposed the LitCab framework, in which a single linear layer is trained over the LLM’s last hidden layer representations to predict a bias term, and the model’s logits are then altered accordingly to update the response confidence. Since these methods are model-specific and need additional task-specific training labels, they cannot be readily applied to new models and tasks. In comparison, as an unsupervised method, semantic density is "off-the-shelf" for any new models and tasks, without need of additional data or modifications to the original LLMs.
A common limitation of all the above methods is that they only consider the lexical information, and do not take into account the semantic relationships between responses. Yet semantics is critical in analyzing LLM outputs. As a fourth direction, semantic entropy [16] is a SOTA technique that quantifies semantic uncertainty for LLMs. It works by grouping the generated samples based on their semantic equivalence, then generating an entropy-based indicator as an uncertainty metric. Although its performances is promising compared to the other approaches, as discussed in Section 1, it has two intrinsic limitations: (1) The generated semantic entropy is prompt-wise instead of response-wise; and (2) Only a one-cut equivalence relationship is considered during semantic analysis. The proposed semantic density improves over these two aspects by providing a response-specific uncertainty indicator and analyzing the relationship among generated responses in a fine-grained semantic space.
Besides the above four major directions, uncertainty quantification for LLMs has been explored from other angles as well. Lin et al. [20] tested several simple baselines, among which a straightforward measurement of semantic dispersion is robust in evaluating selective response generation. Similarly, Manakul et al. [24] proposed a framework with several variants that use sampling consistency for detecting hallucinations. These two studies assume a very restricted condition in which the original sampling likelihoods of each response are not used; thus, the uncertainty information extracted by these methods is limited. The study by Ling et al. [21] focused on a specific in-context learning setting, aiming to decompose the uncertainty of LLMs into that caused by demonstration quality and that caused by model configuration. These experiments were limited to classification problems. Similarly, Hou et al. [10] decomposed the uncertainty into data uncertainty and model uncertainty in a prompt-wise approach. Xiao and Wang [44] studied the connections between hallucination and predictive uncertainty, showing that higher uncertainty is positively correlated with higher chance to hallucinate. This study validates the importance of a reliable uncertainty measurement in detecting hallucinations of LLMs. Finally, Huang et al. [11] perform an explorative study using simple baselines on uncertainty measurement for LLMs, highlighting the need for more advanced uncertainty quantification methods developed exclusively for LLMs. This is the goal for the current paper as well.
3 Methodology
This section first defines the LLM uncertainty quantification problem, then describes the design principles and technical details of each algorithmic component, and concludes with a summary of the entire framework.
3.1 Problem Statement
Given a pre-trained LLM, an input prompt $\bm{x}$, and an output sequence $\bm{y}=[y_{1},y_{2},\cdots,y_{L}]$, where $L$ is the number of tokens in $\bm{y}$, the target is to produce an uncertainty metric that has a monotonic relationship with the probability of $\bm{y}$ to be true. Note that this metric should be response-wise, i.e., it is calculated for a specific $\bm{y}$ given $\bm{x}$. The metric can be used as a quantitative indicator for whether a specific response $\bm{y}$ can be trusted.
3.2 Semantic Space
Theoretically, a semantic space can be any metric space such that a distance function is properly defined to measure the semantic similarity between any two output responses, given the input prompt. Note that such a space is prompt-specific, i.e., each prompt results in a specific semantic space in which the distance function measures the contextual semantic similarity between two responses, treating the prompt as a common context.
More concretely, an oracle semantic space $\mathbb{S}$ is assumed to be a Euclidean space where each point is a $D$-dimensional vector that represents a contextual embedding of response $\bm{y}$ given prompt $\bm{x}$:
$\bm{v}=E(\bm{y}|\bm{x}),$ (1)
where $\bm{v}\in\mathbb{R}^{D}$, and $E(\cdot|\cdot)$ is an encoder that generates text embeddings with the following properties:
  1. 1.
    All the generated embedding vectors are normalized to have a norm of $\frac{1}{2}$:
    $||\bm{v}||=\frac{1}{2},\ \mathrm{for}\ \bm{v}=E(\bm{y}|\bm{x}),\ \forall\bm{x},\bm{y}.$ (2)
    Whereas most existing text embedding models normalize the output vectors to have a norm of 1 [26], they are rescaled to $\frac{1}{2}$ without changing their direction to make it simpler to integrate them into the kernel function (as explained in Section 3.4).
  2. 2.
    Given a prompt $\bm{x}$ and two resulting responses $\bm{y}_{i}$ and $\bm{y}_{j}$, with $\bm{v}_{i}=E(\bm{y}_{i}|\bm{x})$ and $\bm{v}_{j}=E(\bm{y}_{j}|\bm{x})$, the following constraints exist for three extreme cases:
    $||\bm{v}_{i}-\bm{v}_{j}||=\begin{cases}0,\ \mathrm{if}\ \bm{y}_{i}\ \mathrm{and}\ \bm{y}_{j}\ \mathrm{are\ semantically\ equivalent\ given\ context}\ \bm{x}\\ \frac{\sqrt{2}}{2},\ \mathrm{if}\ \bm{y}_{i}\ \mathrm{and}\ \bm{y}_{j}\ \mathrm{are\ semantically\ irrelevant\ given\ context}\ \bm{x}\\ 1,\ \mathrm{if}\ \bm{y}_{i}\ \mathrm{and}\ \bm{y}_{j}\ \mathrm{are\ semantically\ contradictory\ given\ context}\ \bm{x}.\end{cases}$ (3)
    Given the norm requirement in Eq. 2, the above three cases also correspond to $\bm{v}_{i}=\bm{v}_{j}$, $\bm{v}_{i}\perp\bm{v}_{j}$ and $\bm{v}_{i}=-\bm{v}_{j}$, respectively. Note that $||\bm{v}_{i}-\bm{v}_{j}||$ is not restricted to the above three values. It can be any value within $[0,1]$, depending on the semantic similarity between $\bm{y}_{i}$ and $\bm{y}_{j}$ given $\bm{x}$.
  3. 3.
    Given a prompt $\bm{x}$ and three resulting responses $\bm{y}_{i}$, $\bm{y}_{j}$ and $\bm{y}_{k}$, with $\bm{v}_{i}=E(\bm{y}_{i}|\bm{x})$, $\bm{v}_{j}=E(\bm{y}_{j}|\bm{x})$ and $\bm{v}_{k}=E(\bm{y}_{k}|\bm{x})$,
    $||\bm{v}_{i}-\bm{v}_{j}||<||\bm{v}_{i}-\bm{v}_{k}||,\ \mathrm{if}\ \bm{y}_{i}\ \mathrm{is\ semantically\ closer\ to}\ \bm{y}_{j}\ \mathrm{than\ to}\ \bm{y}_{k},\ \mathrm{given\ context}\ \bm{x}.$ (4)
3.3 Semantic Density Estimator
Given the semantic space $\mathbb{S}$ defined in Section 3.2, the underlying probability distribution from which the LLM samples in $\mathbb{S}$ provides critical uncertainty information: If a response is semantically close to many highly probable samples, it should be more trustworthy compared to a response that is semantically distant from the major sampling possibilities. A classical technique for estimating probability density is kernel density estimation (KDE) [32, 28]. However, the standard KDE only works for continuous variables, whereas the LLM outputs are discrete, i.e., sequences of tokens selected from a finite vocabulary. One possible way to extend KDE to accomodate LLM outputs is to build a density estimator as
$\hat{p}(\bm{y}_{*}|\bm{x})=\sum_{i=1}^{M}f_{i}K(\bm{v}_{*}-\bm{v}_{i})=\frac{1}{\sum_{i=1}^{M}n_{i}}\sum_{i=1}^{M}n_{i}K(\bm{v}_{*}-\bm{v}_{i}),$ (5)
where $\bm{x}$ is the input prompt, $\bm{y}_{*}$ is the target response, and $\bm{v}_{*}=E(\bm{y}_{*}|\bm{x})$. In total, $\sum_{i=1}^{M}n_{i}$ reference responses are sampled to facilitate the density estimation, where $M$ is the number of unique samples. Each $\bm{y}_{i}$ represents a unique sample; $n_{i}$ is the number of occurrences of $\bm{y}_{i}$ during sampling, $f_{i}=\frac{n_{i}}{\sum_{i=1}^{M}n_{i}}$ is the relative frequency of $\bm{y}_{i}$ during sampling, and $K(\cdot)$ is a kernel function.
The design of Eq. 5 is similar to an early variant of KDE [29] that was used to handle integer data. However, it has the drawback that it incorporates no knowledge about the sampling probabilities for each $\bm{y}_{i}$. It thus requires a large number of samplings, including a sufficient number of duplicated results, to obtain the relative frequency as an empirical approximation of the sampling probability. This cost becomes prohibitive for LLMs given how expensive LLM inference generally is.
In contrast with the inherently unknown probability distributions in standard KDE, the output token probabilities can be explicitly calculated in LLM sampling, and with this information, a more sample-efficient estimator can be developed. Given a prompt $\bm{x}$ and a resulting response $\bm{y}_{*}$, the semantic density of $\bm{y}_{*}$ is defined as
$\mathrm{SD}(\bm{y}_{*}|\bm{x})=\frac{1}{\sum_{i=1}^{M}p(\bm{y}_{i}|\bm{x})}\sum_{i=1}^{M}p(\bm{y}_{i}|\bm{x})K(\bm{v}_{*}-\bm{v}_{i}),$ (6)
where $\bm{v}_{*}=E(\bm{y}_{*}|\bm{x})$, and $\bm{v}_{i}=E(\bm{y}_{i}|\bm{x})$ for $i=1,2,\cdots,M$. The $M$ unique responses $\bm{y}_{i}$ are the reference responses based on which the semantic density of $\bm{y}_{*}$ is estimated. $K(\cdot)$ is a kernel function which will be specified in Section 3.4, and $p(\bm{y}_{i}|\bm{x})$ (for $i=1,2,\cdots,M$) is the probability for the original LLM to generate sequence $\bm{y}_{i}$ given $\bm{x}$. That is, $p(\bm{y}_{i}|\bm{x})=\prod_{j=1}^{L_{i}}p(y_{i,j}|y_{i,1},y_{i,2},\cdots,y_{i,j-1},\bm{x})$, where $L_{i}$ is the number of tokens in $\bm{y}_{i}$ and $p(y_{i,j}|\cdot)$ is the conditional probability to generate token $y_{i,j}$. Note that in cases where $p(\bm{y}_{*}|\bm{x})$ is available, $\bm{y}_{*}$ can also be used as one of the $M$ reference responses.
One advantage of the semantic density estimator of Eq. 6 is that each result $\bm{y}_{i}$ only needs to be sampled once; their relative frequency $f_{i}$ can then be estimated as $f_{i}=\frac{p(\bm{y}_{i}|\bm{x})}{\sum_{i=1}^{M}p(\bm{y}_{i}|\bm{x})}$. Given a sampling budget of $M$ reference responses, it is therefore desirable that these $M$ samples are unique (duplications will be removed before calculating Eq. 6) and have high sampling probabilities, so that they can cover more sampling regions in the semantic space. In the current implementation, diverse beam search [38], which tends to generate diverse and high probable responses, is used to sample the $M$ unique reference responses.
In practice, length-normalized probability [25, 23] is usually used to correct the length bias in sequence probability. Moreover, temperatures scaling [8, 5, 43] is a simple yet effective method for calibrating the token probabilities during sampling. Both methods can be seamlessly integrated into semantic density: The $p(\bm{y}_{i}|\bm{x})$ in Eq. 6 can be replaced with $L_{i}]{p(\bm{y}_{i}|\bm{x})}$, and the temperature changed during sampling to calibrate each $p(y_{i,j}|\cdot)$.
3.4 Dimension-invariant Kernel
In a standard KDE setup, a commonly used kernel for multi-variate cases is the Epanechnikov kernel [6, 7], which was proved to be the most efficient in terms of asymptotic mean integrated squared error [39]. Its original form is
$K(\bm{v})=\frac{\Gamma(2+\frac{D}{2})}{\pi^{\frac{D}{2}}}(1-||\bm{v}||^{2})\bm{1}_{||\bm{v}||\leq 1},$ (7)
where $D$ is the dimension of vector $\bm{v}$, $\Gamma(\cdot)$ is the gamma function, $||\bm{v}||$ is the 2-norm of $\bm{v}$, and $\bm{1}_{\mathrm{condition}}$ equals 1 if the condition is true, 0 otherwise.
In the semantic density estimator use case, one drawback of the original Epanechnikov kernel is that the normalization coefficient $\frac{\Gamma(2+\frac{D}{2})}{\pi^{\frac{D}{2}}}$ changes with the dimension $D$ of $\bm{v}$. As a result, semantic densities calculated using embeddings with different dimensionalities are incomparable. This issue may limit the flexibility in selecting embedding methodologies for semantic density calculation. However, the normalization coefficient can be removed to make the kernel function simpler and more flexible, without affecting the performance of uncertainty quantification. The kernel function of Eq. 6 thus becomes
$K(\bm{v}_{*}-\bm{v}_{i})=(1-||\bm{v}_{*}-\bm{v}_{i}||^{2})\bm{1}_{||\bm{v}_{*}-\bm{v}_{i}||\leq 1}.$ (8)
Although the resulting kernel function does not meet the normalization requirement in standard KDE, it fits uncertainty quantification well. As long as the norm requirements in Eq. 2,  3 and  4 are fulfilled, any embedding models can be used to generate $\bm{v}$, regardless of the embedding dimensionalities. The outcome of kernel function is always within $[0,1]$, and a kernel value of 1, $\frac{1}{2}$, and 0 correspond to semantically equivalent, irrelevant and contradictory responses, respectively. As a result, the semantic density is also within $[0,1]$, with 1 as the highest semantic density a response can obtain, indicating that all the reference responses are semantically equivalent to it; analogously, it obtains a semantic density of 0 when all reference responses are semantically contradictory to it. This consistency in the value range makes practical applications of semantic density convenient: Practitioners can set a fixed threshold on semantic density to detect unreliable responses.
3.5 Semantic Distance Measurement via the Natural Language Inference (NLI) Model
Although most of the existing text-embedding models work in the semantic space defined in Section 3.2, they do not perform well in measuring semantic similarities [26]. Moreover, they can only consider input texts as a whole instead of doing a contextual inference on part of the input, i.e., they can only obtain $E(\bm{x}+\bm{y})$ instead of $E(\bm{y}|\bm{x})$, where $\bm{x}+\bm{y}$ means a concatenation of $\bm{x}$ and $\bm{y}$.
The natural language inference (NLI) classification model[9] has proven to be effective in analyzing the semantic relationship between LLM responses with the prompt as context [16]. Given a pair of texts, a NLI model performs a classification task and outputs the probabilities for them to be semantically equivalent ("entailment" class), irrelevant ("neutral" class), or contradictory ("contradiction" class). Given the output class probabilities, the expectation of $||\bm{v}_{*}-\bm{v}_{i}||$ can be obtained as
${\begin{array}[]{ll}\mathbb{E}(||\bm{v}_{*}-\bm{v}_{i}||^{2})&=1^{2}\cdot p_{\mathrm{c}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})+(\frac{\sqrt{2}}{2})^{2}\cdot p_{\mathrm{n}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})+0^{2}\cdot p_{\mathrm{e}}(\bm{y}_{*},\bm{y}_{i}|\bm{x}),\\ &=p_{\mathrm{c}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})+\frac{1}{2}\cdot p_{\mathrm{n}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})\end{array}}$ (9)
where $p_{\mathrm{c}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$, $p_{\mathrm{n}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ and $p_{\mathrm{e}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ are the probabilities for $\bm{y}_{*}$ and $\bm{y}_{i}$ to be semantically contradictory ("c" for "contradiction" class), irrelevant ("n" for "neutral" class) and equivalent ("e" for "entailment" class), respectively, given context $\bm{x}$. The expected value of $||\bm{v}_{*}-\bm{v}_{i}||^{2}$ can then be used in Eq. 8 to obtain the kernel function output.
3.6 Summary of the Semantic Density Framework
Algorithm 1 describes how the semantic density metric is deployed on a given task and model. The procedure consists of four main steps, i.e., sampling the reference responses, analyzing semantic relationships, calculating the kernel function, and calculating the semantic density.
Algorithm 1 Procedure for deploying semantic density
0:   $\bm{y}_{*}$: target response that needs uncertainty estimation $\bm{x}$:   original prompt for generating $\bm{y}_{*}$ $M$: number of unique reference responses to be sampled given $\bm{x}$
0:   $\mathrm{SD}(\bm{y}_{*}|\bm{x})$: semantic density for $\bm{y}_{*}$ given $\bm{x}$ Step 1: Reference Response Sampling:
1:  sample $M$ unique reference responses $\bm{y}_{i}$ (for $i=1,2,\cdots,M$) with prompt $\bm{x}$ on the original LLM using diverse beam search, and record each corresponding length-normalized sampling probability $L_{i}]{p(\bm{y}_{i}|\bm{x})}$ Step 2: Semantic Relationship Analysis:
2:  for $i=1\ \mathrm{to}\ M$ do
3:     obtain $p_{\mathrm{c}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ and $p_{\mathrm{n}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ using NLI classification model
4:     calculate expectation $\mathbb{E}(||\bm{v}_{*}-\bm{v}_{i}||^{2})=p_{\mathrm{c}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})+\frac{1}{2}\cdot p_{\mathrm{n}}(\bm{y}_{*},\bm{y}_{i}|\bm{x})$ Step 3: Kernel Function Calculation:
5:  for $i=1\ \mathrm{to}\ M$ do
6:     calculate kernel function value using the expectation of $||\bm{v}_{*}-\bm{v}_{i}||^{2}$, given by:$K(\bm{v}_{*}-\bm{v}_{i})=(1-\mathbb{E}(||\bm{v}_{*}-\bm{v}_{i}||^{2}))\bm{1}_{\mathbb{E}(||\bm{v}_{*}-\bm{v}_{i}||)\leq 1}$ Step 4: Semantic Density Calculation:
7:  calculate semantic density: $\mathrm{SD}(\bm{y}_{*}|\bm{x})=\frac{1}{\sum_{i=1}^{M}\sqrt[L_{i}]{p(\bm{y}_{i}|\bm{x})}}\sum_{i=1}^{M}\sqrt[L_{i}]{p(\bm{y}_{i}|\bm{x})}K(\bm{v}_{*}-\bm{v}_{i})$
Computational Cost:
In terms of computational cost, only the first two steps involve model inferences. The first step utilizes diverse beam search, in which the group number equals $M$ with one beam in each group, and thus only $M$ inferences need to be done by the original LLM. The second step requires another $M$ or $2M$ inferences by the NLI classification model, depending on whether the relationship analysis is performed in a bi-directional manner. Considering the fact that NLI models are usually significantly smaller than LLMs (e.g., the Deberta-large-mnli model [9] used in the implementation in this paper only has 1.5 billion parameters), the computational cost is therefore mainly determined by the LLM inferences in the first step.
4 Experiments
This section first evaluates the performance of semantic density by comparing it with six existing uncertainty metrics over various LLMs and benchmarks. After that, two empirical studies are performed to investigate the robustness of semantic density when the number of reference responses and sampling strategy for target response are varied.
4.1 Performance Evaluation
Table 1: Performance of different uncertainty metrics across various LLMs and datasets
CoQA
AUROC SD SE[16] P(True)[15] Deg[20] NL[25] NE[23] PE[15]
Llama-2-13B 0.783 0.633 0.594 0.734 0.709 0.629 0.647
Llama-2-70B 0.783 0.621 0.576 0.721 0.716 0.617 0.647
Llama-3-8B 0.738 0.599 0.593 0.795 0.676 0.608 0.604
Llama-3-70B 0.789 0.608 0.670 0.729 0.698 0.587 0.641
Mistral-7B 0.788 0.627 0.667 0.737 0.704 0.614 0.632
Mixtral-8x7B 0.786 0.626 0.589 0.728 0.708 0.617 0.651
Mixtral-8x22B 0.791 0.614 0.614 0.726 0.700 0.604 0.649
TriviaQA
AUROC SD SE P(True) Deg NL NE PE
Llama-2-13B 0.848 0.672 0.589 0.824 0.675 0.574 0.556
Llama-2-70B 0.829 0.677 0.556 0.787 0.714 0.582 0.566
Llama-3-8B 0.866 0.662 0.647 0.796 0.834 0.636 0.622
Llama-3-70B 0.828 0.663 0.654 0.764 0.828 0.611 0.596
Mistral-7B 0.866 0.690 0.589 0.828 0.745 0.615 0.536
Mixtral-8x7B 0.846 0.685 0.562 0.797 0.795 0.644 0.605
Mixtral-8x22B 0.829 0.686 0.604 0.762 0.801 0.644 0.607
SciQ
AUROC SD SE P(True) Deg NL NE PE
Llama-2-13B 0.757 0.570 0.572 0.727 0.693 0.513 0.574
Llama-2-70B 0.746 0.643 0.584 0.713 0.637 0.554 0.615
Llama-3-8B 0.780 0.611 0.564 0.731 0.686 0.597 0.651
Llama-3-70B 0.771 0.613 0.556 0.706 0.724 0.558 0.520
Mistral-7B 0.771 0.618 0.568 0.736 0.669 0.565 0.528
Mixtral-8x7B 0.773 0.612 0.585 0.716 0.726 0.612 0.658
Mixtral-8x22B 0.775 0.620 0.602 0.719 0.715 0.602 0.628
NQ
AUROC SD SE P(True) Deg NL NE PE
Llama-2-13B 0.689 0.581 0.592 0.686 0.588 0.571 0.640
Llama-2-70B 0.676 0.545 0.531 0.691 0.567 0.573 0.620
Llama-3-8B 0.710 0.583 0.517 0.706 0.601 0.603 0.615
Llama-3-70B 0.723 0.577 0.643 0.714 0.631 0.603 0.615
Mistral-7B 0.680 0.597 0.523 0.676 0.640 0.635 0.631
Mixtral-8x7B 0.729 0.599 0.576 0.720 0.654 0.603 0.608
Mixtral-8x22B 0.709 0.577 0.504 0.704 0.638 0.625 0.680
Following the usual evaluation approach in the literature [16], the uncertainty metric is used as a quantitative indicator of how likely the response is going to be correct. Uncertainty values above a threshold are taken as incorrect while those below are taken as correct. For each threshold, the true positive rate vs. false positive rate is then measured. The area under this curve, namely area under receiver operator characteristic curve (AUROC), is calculated for each uncertainty metric. The AUROC score equals the probability that a randomly chosen incorrect response has a higher uncertainty than a randomly chosen correct response. A perfect uncertainty metric would have an AUROC score of 1 while a random metric would have 0.5. Note also that a higher semantic density corresponds to a lower uncertainty.
The performance of semantic density (SD) was compared with six existing LLM uncertainty quantification methods: semantic entropy (SE) [16], P(True) [15], degree (Deg) [20], length-normalized likelihood (NL) [25], length-normalized entropy (NE) [23] and predictive entropy (PE) [15]. These methods were applied to seven SOTA open-source LLMs: Llama-2-13B [37], Llama-2-70B [37], Llama-3-8B [2], Llama-3-70B [2], Mistral-7B [12], Mixtral-8x7B [13] and Mixtral-8x22B [35]. Each LLM was tested on four free-form question-answering datasets commonly used in the literature: CoQA [30], TriviaQA [14], SciQ [40] and Natural Questions (NQ) [17]. For each question, 10 responses were generated using group beam search and used as reference responses in calculating SD, SE, Deg, NE and PE (note that P(True) and NL do not need reference responses). Each unique response among these 10 will also be used as a target response, e.g., the response that needs an uncertainty estimation, in calculating the AUROC scores of uncertainty metrics. Detailed experimental configuration and parametric setup is described in Appendix A.1.
Table 1 shows the AUROC scores of each uncertainty metric across different models and datasets, with the best entry in each configuration highlighted in boldface. SD performs best in 26 out of 28 cases, demonstrating that it is reliable and robust as an uncertainty metric for LLM responses. In two cases it is outperformed by Deg. After investigation, the inherent sequence likelihood returned by the original LLM was badly calibrated in these two cases. Deg is the only method that ignores the likelihood information during its calculation, making its performance unaffected by this negative factor. However, for the other 26 cases, SD is able to utilize the likelihood information to its advantage, and outperform Deg.
To confirm that the observed differences are statistically significant, a paired $t$-test (paired by LLM and dataset) was performed between SD and the other metrics (Table 2 in Appendix A.2). The $p$-values are consistently below $10^{-6}$, indicating that the performance gains of SD are strongly statistically significant.
4.2 Robustness of Semantic Density
Two additional empirical studies were performed to evaluate the robustness of semantic density when the number of reference responses varies or the sampling strategy for target response changes.
In the first study, the number of reference responses was reduced from 10, which is a standard setup for existing methods [16], to one, which is the extreme minimum case. Figure 1 shows the resulting AUROC scores, covering the same four datasets and seven LLMs. Although performance indeed decreases with fewer reference responses, the decrease is minor as long as the number of references is at least four. This result suggests that semantic density can provide reasonable performance even with a very limited budget for reference sampling.
Figure 1: AUROC scores of semantic density with one to 10 reference responses. Each subfigure corresponds to the dataset indicated by the subfigure heading. Each curve corresponds to the base LLM identified in the legend below the subfigures. Providing more reference responses generally increases the reliability of semantic density, but in most cases only four samples are sufficient.
In real-world applications, users may have different preferences when using LLMs to generate responses: Some may prefer a greedy sampling strategy while others may need diverse responses. The second study thus investigated how each uncertainty metric performs when the target response is sampled using different such strategies. The diverse beam search method inherently utilizes different strategies for each beam group: The first group performs a greedy beam search while later groups encourage more diverse responses. Following the setup in Section 4.1, the AUROC scores were calculated for target responses from each group separately, and the results averaged over the four datasets.
As the results in Figure 2 show, semantic density exhibits consistently good AUROC scores across different beam groups. Thus, it is robust against both more greedy and more diverse sampling strategies, thus covering a range of possible user preferences. In contrast, other approaches either perform consistently worse compared to semantic density across different beam groups, or their performance is unstable when the sampling strategy changes.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: AUROC scores over the different beam groups. The plots show the average AUROC scores over the four datasets for the different beam groups in diverse beam search. Smaller group index corresponds to the group with more greedy generation strategy, while the group with larger index tends to be more diverse during response generation. Each subfigure corresponds to one of the seven LLMs, as indicated by the subfigure heading. Each curve represents one uncertainty metric indicated by the legend at lower right. Semantic density exhibits consistently better and stable performance across different groups, compared to other methods.
5 Discussion and Future work
In terms of the broader societal impact of this work, the proposed semantic density provides a general way to evaluate the trustworthiness of responses generated by LLMs. This ability should have a positive impact on real-world applications that are safety-critical, such as healthcare and finance. Practitioners can utilize semantic density as an off-the-shelf indicator to filter out unreliable responses.
One limitation of semantic density is that it needs access to the output probabilities of generated tokens, which may not be available in some proprietary LLMs. In such a case, the more expensive variant in Eq. 5 can be considered as an alternative. Since semantic density does not require any further access to the internal states or weights of the original LLMs, it is still widely applicable.
The framework for measuring semantic density is modular, and therefore its main components can be extended in future work. First, new strategies that explicitly encourage a better coverage of semantic space can be developed to generate reference responses. Such extensions should improve the reliability of semantic density further. Second, text embedding methods that can measure contextual semantic similarity between responses more reliably will be helpful as well. Third, kernel functions specifically designed for the semantic space should allow utilizing semantic relationships more efficiently. Fourth, more precise methods for calibrating inherent token probabilities will form a more reliable base for calculating semantic density.
6 Conclusion
This paper proposes semantic density as a practical new metric for measuring uncertainty in LLM responses. It overcomes the limitations of existing approaches by utilizing the semantic information in an efficient and precise manner. It is response-specific, "off-the-shelf", and applicable to free-form generation tasks. Experimental comparisons with six existing uncertainty metrics across seven SOTA LLMs and four benchmark datasets suggest that it is accurate, robust, and general, and can therefore help deploy LLMs in safety-critical domains.
Appendix A Appendix
A.1 Experimental Setup
This section provides the detailes of experimental setup for reproducing the results presented in Section 4.
Base LLMs:
For all the seven base LLMs, the open-source versions in Huggingface Transformers library [41] were used in the experiments. More specifically, the following versions are used: meta-llama/Llama-2-13b-hf, meta-llama/Llama-2-70b-hf, meta-llama/Meta-Llama-3-8B, meta-llama/Meta-Llama-3-70B, mistralai/Mistral-7B
-v0.1
, mistralai/Mixtral-8x7B-v0.1, and mistralai/Mixtral-8x22B-v0.1. When running diverse beam search in generating the responses, the generate() function is used with diversity_penalty=1.0, a default temperature of 1.0, and the num_beam_groups equals the number of beams so that each group has exactly one beam.
Datasets:
Following the setup in Kuhn et al. [16], for all the datasets, an answer is considered to be correct if its Rouge-L [18] to any of the reference answers is larger than 0.3, after trimming the redundant continuations. The details of the datasets are as follows:
  • CoQA: The coqa-dev-v1.0 version is used with 1596 questions randomly selected for the experiments, using the Huggingface datasets.train_test_split function with a seed value of 10. The prompt format follows the same setup as in Kuhn et al. [16].
  • TriviaQA: The dataset is loaded with the Huggingface datasets API. 1705 questions were randomly selected from the validation split for the experiments, using the Huggingface datasets.train_test_split function with a seed value of 10. A 10-shot prompt format is used following the setup in Kuhn et al. [16].
  • SciQ: The test split from https://github.com/launchnlp/LitCab is used in the experiments. It contains 990 questions. A 10-shot prompt format is used following the setup in Kuhn et al. [16].
  • NQ: The test split from https://github.com/launchnlp/LitCab is used in the experiments. 1800 questions were randomly selected using the Huggingface datasets.train_test_
    split
    function with a seed value of 10. A 10-shot prompt format is used following the setup in Kuhn et al. [16].
Uncertainty Metrics:
The exactly same target response and reference responses are used for all the tested methods. For SD, SE and Deg, the microsoft/deberta-large-mnli model from Huggingface Transformers library [41] is used as the NLI classification model. Following Kuhn et al. [16], the probabilities for "contradiction", "neutral" and "entailment" are averaged bidirectionally. The parametric setup for the uncertainty metrics tested in Section 4 is summarized as follows:
  • semantic density (SD): The exactly same steps as described by Algorithm 1 were implemented, with a fixed temperature of 0.1 applied to rescale the value of each token probability during postprocessing.
  • semantic entropy (SE): The original implementation and parametric setup from https://github.com/lorenzkuhn/semantic_uncertainty was used.
  • P(True): The original few-shot prompt format from Kadavath et al. [15] is used.
  • degree (Deg): The "entailment" probability returned by the NLI model (averaged bidirectionally) is used as the similarity between two responses, which is then used as the diagonal element in degree matrix.
  • length-normalized likelihood (NL): The original form as in Murray and Chiang [25] was implemented.
  • length-normalized entropy (NE): The implementation from https://github.com/lorenzkuhn
    /semantic_uncertainty
    was used.
  • predictive entropy (PE): The implementation from https://github.com/lorenzkuhn/semantic
    _uncertainty
    was used.
Compute Resources:
All the experiments in Section 4 were running on an AWS P4de instance with 96 Intel(R) Xeon(R) Platinum 8275CL CPU @ 3.00GHz, 1152GB memory and 8 NVIDIA A100 (80GB). The GPU memory required to run the experiments is dependent on the base LLMs, as detailed below:
  • Llama-2-13B: $\sim$30GB GPU memory.
  • Llama-2-70B: $\sim$140GB GPU memory.
  • Llama-3-8B: $\sim$20GB GPU memory.
  • Llama-3-70B: $\sim$140GB GPU memory.
  • Mistral-7B: $\sim$20GB GPU memory.
  • Mixtral-8x7B: $\sim$110GB GPU memory.
  • Mixtral-8x22B: $\sim$300GB GPU memory.
The exact computation time was affected by many factors, e.g., the current workload of the machine, the prompt length of the question, the generated response length, whether the cache option is turned on to store the model states of LLMs, etc.
A.2 Results of Statistical Tests
Table 2 shows the $p$-values of the paired $t$-tests as described in Section 4.1.
Table 2: Statistical significance of SD’s advantage over other methods ($p$-values of paired $t$-tests)
SD vs. SE P(True) Deg NL NE PE
4.83E-17 1.71E-15 1.16E-7 4.62E-14 2.35E-15 4.63E-13